OpenStudy (anonymous):

Find F'(x) when F(x)=sin(-x^2) intergral from 0 to x^2 of e^(t)^(2)dt.

zepdrix (zepdrix):

Ooo interesting problem! :)

zepdrix (zepdrix):

\[\Large\rm F(x)=\sin(-x^2)\int\limits_0^{x^2} e^{t^{2}}~dt\] Find \(\Large\rm F'(x)\). This is the problem? That look correct?

OpenStudy (anonymous):

Yes.

zepdrix (zepdrix):

I'm gonna use some shorthand notation for a sec, \[\Large\rm F(x)=\sin(-x^2)\cdot Int\]So to find the derivative, we apply product rule, yes?\[\Large\rm F'(x)=\color{royalblue}{\left[\sin(-x^2)\right]'}\cdot Int+\sin(-x^2)\cdot \color{royalblue}{\left[Int\right]'}\]

zepdrix (zepdrix):

So we need the derivative of these blue parts. Understand how to deal with the first one?

OpenStudy (anonymous):

Okay, but where do I go from there? So far in calculus we have only learned the fundamental theorem of calculus. So they expect us to do it that way.

zepdrix (zepdrix):

You apply the Fundamental Theorem of Calculus, Part 1, to take the derivative of that integral. We can go over that in a sec if you need. But first deal with the sin(-x^2) derivative. You just apply normal rules to deal with that. Do you remember how?

OpenStudy (anonymous):

Yeah, that would be -2cos(x^2), right?

zepdrix (zepdrix):

\[\Large\rm (\sin(-x^2))'=\cos(-x^2)(-2x)\]If you want to absorb the negative inside the cosine function, (since it's an odd function) I guess you can do that...\[\Large\rm =-2x \cos(x^2)\] but it's probably better to leave the negative inside alone.\[\Large\rm -2x \cos(-x^2)\]

zepdrix (zepdrix):

Yours looks a little different, do you understand where the x came from? Derivative of x^2 is 2x, right?

OpenStudy (anonymous):

Yeah, I just wrote it down wrong when taking the derivative. Sorry about that. I just redid it and got the same answer as you.

zepdrix (zepdrix):

cool c: so for the derivative of the integral, it's a little weird! We'll apply the FTC, part 1:\[\Large\rm \frac{d}{dx}\int\limits_0^x f(t)~dt=f(x)\]

OpenStudy (anonymous):

So that would be \[e ^{x ^{2}}\].

zepdrix (zepdrix):

It's a bit more complicated than that since our upper limit isn't just "x". Lemme try to explain this in a way that will make sense...

OpenStudy (anonymous):

Okay.

zepdrix (zepdrix):

\[\huge\rm \int\limits\limits_0^{x^2} e^{\color{orangered}{t}^{2}}~dt\quad=\quad \int\limits\limits_0^{x^2} f(\color{orangered}{t})~dt\]So you correctly identified the f(t) that we're integrating.

zepdrix (zepdrix):

Integrating will give us "something", let's call it F(t) the anti-derivative. This is the part that is a little hard for some students to understand: We're not actually looking for an anti-derivative, we're assuming one exists and writing it as F(t). We'll plug in our boundaries:\[\huge\rm F(\color{orangered}{t})|_0^{x^2}\quad=\quad F(\color{orangered}{x^2})-F(\color{orangered}{0})\]

zepdrix (zepdrix):

Then if we want to take the derivative, we end up back at our original f function,\[\huge f(\color{orangered}{x^2})\cdot 2x-0\]

zepdrix (zepdrix):

Since the inner function (x^2) was more than just "x", I had to apply the chain rule.

zepdrix (zepdrix):

I know I know it's a weird process!! ~You anti-differentiate ~plug in boundaries ~differentiate So you end up with the same thing you started with, but it's got a little baggage along with it now.

OpenStudy (anonymous):

Isn't that the second step of the fundamental theorem of calculus though?

zepdrix (zepdrix):

True I guess it is! :) I'm using both steps of the fundamental theorem to show what's going on.

OpenStudy (anonymous):

Oh, okay!

zepdrix (zepdrix):

F(0) is some anti-derivative of e^{t^2} with 0 plugged in. So the whole thing ends up being a costant value. When you take the derivative of a constant, you get zero right? That's why F(0) is giving us zero when we take it's derivative with respect to x.

zepdrix (zepdrix):

\[\huge =f(\color{orangered}{x^2})\cdot 2x\]\[\huge\rm =e^{(\color{orangered}{x^2})^2}\cdot 2x\]

zepdrix (zepdrix):

You have to plug the entire x^2 into the t. See how we get an extra square because of the square that was already on the t?

OpenStudy (anonymous):

Oh, that makes a lot more sense now.

zepdrix (zepdrix):

Yah this is a tricky problem! If you have a good grasp on the FTC, Part 2, it really helps here. When you're able to setup your F(b)-F(a), you can see where the pieces are coming from, like that 2x.

OpenStudy (anonymous):

So would you have to do anything else with the derivative of sin(-x^2)?

zepdrix (zepdrix):

cos(-x) = cos(x) because cosine is an even function* blah I said odd before.. mistake. So you can write your \[\Large\rm \cos(-x^2)\]as\[\Large\rm \cos(x^2)\]if you want. But that doesn't seem necessary.

zepdrix (zepdrix):

\[\Large\rm F'(x)=\color{royalblue}{\left[\sin(-x^2)\right]'}\cdot Int+\sin(-x^2)\cdot \color{royalblue}{\left[Int\right]'}\] \[\Large\rm F'(x)=-2x\cos(-x^2)\cdot Int+\sin(-x^2)\cdot 2x\cdot e^{x^4}\]Something like that yah? :o

zepdrix (zepdrix):

With the integral part written out* I'm too lazy though lol

OpenStudy (anonymous):

Okay, thank you so much for all your help!